OpenStudy (anonymous):

assume that angle A+angle B=180 degrees. which of the following is equal to cos B A. Sin A B. -Sin A C. Cos B D. - Cos B

OpenStudy (anonymous):

pls EXPLAIN how u get the answer

OpenStudy (anonymous):

Maybe I'm missing something... Isn't cosB equal to cosB?

OpenStudy (anonymous):

The other possible answer is -cosA but that's not there

OpenStudy (raden):

A+B=180 B = 180-A put cos to both sides, giving us cosB = cos(180-A) or cosB = -cosA

OpenStudy (anonymous):

oops sorry the question asked for Cos A

OpenStudy (anonymous):

Well that's solved then :)

OpenStudy (anonymous):

yes! thanks @RadEn :)

OpenStudy (anonymous):

and thanks BluFoot!